You just found out that you have $539.74 in a savings account. Your aunt opened the account in your name exactly 6 years ago, and deposited $X in the account. Then she put in another $58 three years after opening the account. Other than $X and the $58, there were no other deposits. If the interest rate for the past 6 years has been 4%, what is X? (choose the closest number) $396 $386 $375 $351
Answer #1

Answer:Option C: $375

Amount in my account=$539.74

interest rate r=4%

FV of X=X*(1+r)^6=X*(1+4%)^6=1.265X

FV of $58=58*(1+r)^3=58*(1+4%)^3=$65.24

FV of X + FV of 58=539.74

1.265X+65.24=539.74

X=$375